OpenStudy (anonymous):

Write the equation of the line in slope-intercept form that passes through the point (3, -2) and has a slope of 2.

OpenStudy (abb0t):

\[(Y-Y_1)=m(X-X_1)\] where \[Y_1 \] is ANY y-value (if you're given two (x,y) values. and \[x_1 \] is also ANY x value. but in ur case, u only are given one (x,y) so just plug it in and rearrange to make it look like y= mx +b where m= slope.

OpenStudy (anonymous):

y = 2x + 7

OpenStudy (anonymous):

@abb0t

OpenStudy (anonymous):

@en28so

OpenStudy (abb0t):

yes

OpenStudy (anonymous):

is that it